Mickey, a daredevil mouse of mass 0.0231 kg, is attempting to become the world's first "mouse cannonball." He is loaded into a spring‑powered gun pointing up at some angle and is shot into the air. The gun's spring has a force constant of 92.3 N/m and is initially compressed a distance of 0.147 m from its relaxed position. If Mickey has a constant horizontal speed of 2.17 m/s while he is flying through the air, how high ℎ above his initial location in the gun does Mickey soar? Assume ?=9.81 m/s2. h=?
Answer #1

Initial energy, E1 = 0.5 x 92.3 x 0.147^2 = 0.997 J

Final energy, E2 = 0.5 x 0.0231 x 2.17^2 = 0.054 J

Using law of conservation of energy,

mgh = 0.997 - 0.054

Height, h = (0.94)/(0.0231 x 9.81) = 4.16 m

Comment in case any doubt please rate my answer....